### v1

# Communication to Completion: Modeling Collaborative Workflows with Intelligent Multi-Agent Communication

Abstract
Teamwork in workspace for complex tasks requires diverse communication strategies, but current multi-agent LLM systems lack systematic frameworks for task oriented communication. We introduce Communication to Completion (C2C), a scalable framework that addresses this gap through two key innovations: (1) the Alignment Factor (AF), a novel metric quantifying agent task alignment that directly impacts work efficiency, and (2) a Sequential Action Framework that integrates stepwise execution with intelligent communication decisions. C2C enables agents to make cost aware communication choices, dynamically improving task understanding through targeted interactions. We evaluated C2C on realistic coding workflows across three complexity tiers and team sizes from 5 to 17 agents, comparing against no communication and fixed steps baselines. The results show that C2C reduces the task completion time by about 40% with acceptable communication costs. The framework completes all tasks successfully in standard configurations and maintains effectiveness at scale. C2C establishes both a theoretical foundation for measuring communication effectiveness in multi-agent systems and a practical framework for complex collaborative tasks.
Communication to Completion: Modeling Collaborative Workflows with Intelligent Multi-Agent Communication
Yiming Lu1,2thanks: Work done during internship at Zoom, Xun Wang2, Simin Ma2, Shujian Liu2, Sathish Reddy Indurthi2 Song Wang2, Haoyun Deng2, Fei Liu1, Kaiqiang Song2 1Emory University  2Zoom Video Communications {yiming.lu,fei.liu}@emory.edu {xun.wang,kaiqiang.song}@zoom.us
1 Introduction
Figure 1: Communication to Completion architecture: a five-layer stack simulation engine
Modern organizations run on teams: people plan, divide, and verify work under deadlines while juggling synchronous meetings, asynchronous chats/emails, and limited attention. Extending this paradigm to AI, coordinating teams of LLM agents to execute complex, long-horizon tasks (e.g., software development) remains a grand challenge (Park et al., 2023; Hong et al., 2024; Qian et al., 2023). Effective collaboration hinges on communication: too much creates coordination overhead; too little yields misalignment and rework.
Recent multi-agent frameworks demonstrate that structured dialogues and role specialization can transform general purpose LLMs into cooperative problem solvers across software engineering, data analysis, and decision support (Wu et al., 2024; Hong et al., 2024; Qian et al., 2023; Li et al., 2023). Beyond mere tool use, conversation itself decomposes tasks, aligns partial knowledge, and arbitrates proposals, often outperforming single agent pipelines on complex problems (de Souza Santos and Ralph, 2022; Khan et al., 2024; Yao et al., 2023a; Besta et al., 2024). Yet many systems still schedule or trigger communication via fixed or purely reactive heuristics that ignore the evolving trade-off between coordination cost and task progress (Wu et al., 2024; Chen et al., 2023).
Meanwhile, empirical studies of software teams warn that coordination costs (especially meetings), can dominate delivery time in modern distributed work, with a post-2020 shift toward heavier synchronous load and changing patterns across chat, email, and meetings (Mok et al., 2023; Yang et al., 2022). These observations motivate treating communication as a first class resource to be scheduled, routed, and evaluated rather than as a byproduct of agent reasoning, thus leading to a central open question: how can agents learn to communicate strategically to maximize collaborative benefit while minimizing coordination costs?
In this work, we address the above question by casting multi-agent collaboration as cost communication and optimizing it accordingly. Our contributions are:
  • We introduce Communication to Completion (C2C). A scalable framework models when, whom, and how to communicate as part of execution and instantiate it with a Sequential Action Framework (SAF) that synchronizes agent actions into discrete timesteps for deterministic, reproducible collaboration.
  • We define Alignment Factor (AF), a metric that quantifies each agent’s task understanding and directly modulates work efficiency. Communications update AF, and AF in turn scales work rate, linking conversation to progress and yielding interpretable traces.
  • We demonstrate through comprehensive experiments spanning complexities and team sizes, C2C lowers completion time and improves work efficiency compared to baselines.
2 Related Work
Multi-Agent Collaboration Frameworks
LLM agents are increasingly organized as teams rather than monoliths. Social simulation work demonstrates emergent, human-like collective behavior (e.g., the “Smallville” agents of Park et al. (2023) and the recent large scale AgentSociety simulation (Piao et al., 2025)). For goal oriented collaboration—particularly software development—frameworks such as MetaGPT (Hong et al., 2024) and ChatDev (Qian et al., 2023) impose predefined roles and SOPs that mirror corporate workflows. AutoGen (Wu et al., 2024) offers a flexible substrate for composing conversable agents, while AgentVerse (Chen et al., 2023) and OpenAgents (Xie et al., 2023) provide extensible environments for building and deploying general purpose agents. Outside pure research prototypes, platforms like OpenHands operationalize multi/single agent development loops in realistic toolchains (Wang et al., 2025). In parallel, role-playing and deliberation structures further organize collaboration within conversations (Li et al., 2023; Yao et al., 2023a; Besta et al., 2024), and new evaluations emphasize interactive, multi-environment testing and repository-level search and planning (Liu et al., 2023a; Antoniades et al., 2024). Our work differs in elevating communication from a procedural step to an optimizable resource, explicitly linking when/what to communicate with collaborative progress via an alignment driven mechanism.
Communication in Multi-Agent Systems
Communication has long been central in multi-agent learning. Classic MARL studies learned protocols end-to-end under partial observability (Foerster et al., 2016; Sukhbaatar et al., 2016). Recent advances focus on efficient or sparse communication to reduce overhead, e.g., decentralized scheduling and model-based message estimation (Han et al., 2023; Liu et al., 2023c; Kim et al., 2019; Singh et al., 2018). In LLM agent settings, natural-language messaging carries high-level, semantic content, and methods increasingly structure agent talk through reasoning–acting loops (ReAct) (Yao et al., 2023b), self improvement with textual feedback (Reflexion) (Shinn et al., 2023), and iterative self-refinement (Madaan et al., 2023). Multi-agent debate further uses conversational arbitration to improve correctness and planning (Du et al., 2023). Yet many systems still trigger communication on fixed or purely reactive heuristics. We instead propose an alignment driven policy that decides when, whom and how to communicate based on a measurable, task grounded alignment signal.
Figure 2: C2C timeline on a software bug-fix task. Step 1: the manager decomposes the task and assigns subtasks. Step 2: workers execute intentions and produce interim results. Communications update the Alignment Factor (AF); higher AF increases work efficiency and reshapes subsequent actions (green feedback). Vertical dashed dividers denote discrete timesteps in the Sequential Action Framework; the cycle repeats until completion (Step N).
Quantifying Coordination and Alignment
Coordination costs are a foundational concern in collaborative systems (Malone, 1988; Brooks Jr, 1995). However, evaluations of LLM multi agents often emphasize final outcomes (e.g., pass rates on benchmarks) over process efficiency. In software engineering, SWE-bench exposes the gap between realistic issue resolution and model capabilities (Jimenez et al., 2023), while SWE-agent (Yang et al., 2024) and OpenHands (Wang et al., 2025) reveal the importance of tool use and orchestration—but still provide limited real-time signals about team alignment. Complementary efforts argue for richer process traces and interactive environments to analyze coordination choices (Liu et al., 2023a; Antoniades et al., 2024), and work on LLM-as-judge warns that automated evaluators may introduce bias at the process level (Liu et al., 2023b; Chen et al., 2024; Tan et al., 2024). We contribute the Alignment Factor (AF) as a dynamic, real-time proxy for shared task understanding that both modulates work efficiency and directly informs agents’ communication decisions, thereby connecting conversational actions to tangible task outcomes.
3 C2C Framework
Communication to Completion is a time-stepped framework for deciding when, whom, and how to communicate in multi-agent collaboration. Figure 2 provides an overview of its mechanism using a software task. We organize the discussion as follows: §3.1 introduces the Sequential Action Framework that ensures deterministic execution, §3.2 describes the Alignment Factor mechanism for quantifying task understanding, §3.3 details how agents make communication decisions, §3.4 explains hierarchical task decomposition and tracking, and §3.5 presents the intention-based agent decision making that integrates these components.
3.1 Sequential Action Framework
Concurrent multi-agent execution often introduces temporal ambiguity. The Sequential Action Framework (SAF) constrains each agent to exactly one action per timestep, yielding deterministic state transitions and reproducible analysis.
Actions
SAF defines four actions capturing common collaborative behavior that capture common collaborative behavior: work (task execution with time cost), communicate (compose/send a message), reply (respond to an incoming message), and meeting (synchronous group discussion). Each action is temporally bounded and commits a single transition.
Synchronized Timesteps
All agents share a fixed temporal grid. At timestep $t$, agent $i$ selects and executes $a_{i}^{t}$, then the system advances to t+1t{+}1. Formally, all actions at $t$ finish before any action at t+1t{+}1 begins:
$\forall i,j:\;a_{i}^{t}\;\text{completes before}\;a_{j}^{t+1}\;\text{starts.}$ (1)
Forward-only Delivery.
Messages sent at $t$ are delivered at t+1t{+}1, preventing instantaneous feedback loops and enforcing causal consistency. A communication buffer stages pending deliveries and flushes at timestep boundaries.
3.2 Alignment Factor
The Alignment Factor (AF) measures an agent’s task-specific understanding and directly modulates effective productivity. Unlike static capability scores, AF evolves through communication. For agent $i$ on task $j$, $AF_{i,j}\in[0.01,1.00]$ denotes comprehension quality. Values near $0.01$ indicate limited understanding; values near $1.00$ indicate mastery. We initialize $AF_{i,j}{=}0.3$ to reflect the initial understanding.
When agent $i$ receives reply about task $j$ from agent $k$, AF is updated based on the communication’s impact on task understanding. An LLM evaluates the message content, agent $i$’s current task context, and the relevance of the received information to determine the new alignment:
$AF_{i,j}^{\text{new}}=\min\!\bigl(1.0,\;AF_{i,j}^{\text{old}}+\Delta_{\text{eval}}\bigr),$ (2)
where $\Delta_{\text{eval}}\in[0,0.5]$ is computed by the LLM based on: (i) how well the message addresses the agent’s knowledge gaps, (ii) the relevance of the information to the specific task requirements, and (iii) the clarity and actionability of the guidance provided. Different message types naturally yield different impacts, help requests addressing critical blockers typically produce larger alignment gains.
If agent $i$ spends $h$ hours on task $j$, the effective progress is
$\text{EffectiveProgress}\;=\;h\cdot AF_{i,j},$ (3)
A low AF impedes an agent’s progress, creating a natural incentive for it to seek clarification and coordination before investing substantial effort.
3.3 Agent Communication Decisions
Agents autonomously decide when, with whom, and how to communicate based on their current task state and situational awareness, without relying on fixed thresholds or rigid scheduling. For clarity, we decompose this decision process into four key components: initiating the communication, selecting the recipients, choosing the appropriate channel, and composing the message content.
Communication Initiation
Agents may initiate communication in response to various task situations: encountering technical difficulties (help_request), facing unclear requirements (need_clarification), identifying coordination needs with team members (meeting_invite), or reaching milestones that warrant status updates (progress_update). The decision to communicate emerges from the agent’s assessment of its current alignment and task complexity.
Recipient Selection
When initiating communication, agents consider multiple factors in selecting recipients: skill relevance to the issue, existing task dependencies, and historical interaction. The selection process reflects realistic collaboration patterns observed in human teams.
Channel Selection
Agents choose among three communication channels. Chat suits quick questions and simple clarifications with rapid response times. Email handles detailed explanations with longer response windows. Meetings address complex coordination needs requiring synchronous discussion among multiple participants.
Message Composition
Communication content is generated based on the agent’s current context, including task description, encountered difficulties, current alignment factor, and relevant technical details. Messages are composed to be informative and actionable, providing recipients with sufficient context to offer effective assistance.
3.4 Hierarchical Task Management
Complex tasks are decomposed by manager agents and tracked as a directed acyclic graph (DAG) of subtasks and dependencies.
The manager performs task decomposition in the beginning of the simulation, by analyzing the task requirements and proposing several subtasks according to team size and skills. Subtasks are connected through dependency edges. The manager tracks these dependencies and coordinates task assignments to ensure workers receive suitable subtasks. And the manager updates the task graph as workers complete subtasks, monitoring overall progress. Parent task progress is computed as a weighted average of subtask progress:
$\text{ParentProgress}=\frac{\sum_{i\in\text{subtasks}}w_{i}\,P_{i}}{\sum_{i\in\text{subtasks}}w_{i}},$ (4)
where $w_{i}$ represents the effort estimate for subtask $i$. The parent task is marked complete only when all required subtasks reach done status, ensuring accurate project tracking.
3.5 Intention-Based Agent Decision Making
Each agent makes decisions through an intention-based mechanism that evaluates the current situation and generates contextually appropriate actions.
Context Formation
At each timestep, the agent constructs a comprehensive context that includes: (i) currently assigned tasks and their completion status, (ii) alignment factors for each assignment indicating task understanding, (iii) recent communications including pending requests and received guidance, and (iv) team state including teammate skills and availability
Intention
Given this context, the agent uses an LLM to carefully reason about the current situation and generate an intention, a high-level decision about what action to take next. The generated intention is then translated into a concrete action within the SAF framework:
  • Work: Continue task execution with effectiveness modulated by current AF.
  • Help: Compose the message identifying difficulties and required expertise (HELP_REQUEST).
  • Clarification: Request additional task details or requirement explanations (NEED_CLARIFICATION).
  • Coordination: Propose synchronous discussion for complex collaboration needs (MEETING_INVITE).
  • Reporting: Share subtasks progress with manager (PROGRESS_REPORT).
When generating communication intentions, the agent also determines appropriate recipients and selects suitable channels (chat, email, meeting) based on urgency and complexity.
Adaptive Behavior
This intention-based method enables agents to exhibit adaptive behavior without hard coded rules. The communication patterns emerge naturally from the agents’ contextual reasoning rather than threshold triggers.
By integrating these components, C2C enables agents to collaborate strategically and adaptively based on actual task needs.
4 Experiments
4.1 Experimental Setup
Metric Complexity No Communication Fixed Steps C2C
Task Completion Rate (%) Simple 100 100 100
Medium 100 100 100
Complex 100 100 100
Avg Completion Time (hours) Simple 7 5 5.5
Medium 20 14.75 13
Complex 33.5 36.25 24.75
Communication Cost Simple 2.03 1.94
Medium 2.75 3.26
Complex 8.12 7.02
Alignment Score (AF) Simple 0.3 0.55 0.51
Medium 0.3 0.59 0.68
Complex 0.3 0.53 0.55
Efficiency Simple 1.14 1.6 1.45
Medium 1.20 1.63 1.85
Complex 1.19 1.10 1.62
Table 1: Comparison of three policies for a 1M+4W team across task complexities (Simple=8h, Medium=24h, Complex=40h). We report Task Completion Rate (%), Avg Completion Time (hours), Communication Cost (total communication time), Alignment Score (mean AF over the run; higher indicates better shared understanding), and Efficiency == completed work / total time. C2C achieves the best time and efficiency on Medium and Complex with comparable communication cost. Bold marks the best value within each row; "-" denotes not applicable.
To evaluate the performance of C2C on realistic tasks, we implement experiments on software engineering workflows across three complexity tiers: Simple (8 hours, basic SWE operations), Medium (24 hours, API integration with authentication), and Complex (40 hours, full stack user authentication system); see Appendix A for details. Each task requires diverse skills including front-end development, back-end systems, database management, security implementation, and testing. Tasks are decomposed into several subtasks according to task complexity and team size with explicit skill requirements and dependencies, mirroring real-world software development scenarios.
Team Configurations
We systematically vary team sizes from 5 to 17 agents: 1M+4W (5 agents), 1M+8W (9 agents), and 1M+16W (17 agents), where M denotes manager and W denotes worker. Additionally, we evaluate multi-task scenarios with 1M+8W handling 2 concurrent tasks to assess workload distribution and task interference effects. Each agent is assigned 2–4 complementary skills from a skill pool, including skills like frontend, backend, and database.
Models and Implementation
All agents are powered by GPT-4o with temperature 0.7 for human like decision making. The Sequential Action Framework operates with 0.25 hour time steps over a maximum of 160 simulation steps (40 hours). Communication costs are calculated using realistic time models: email drafting (9 minutes base + content length), chat messages (3 minutes base), meetings (30 minutes minimum + preparation time + number of participants). We implement thread-depth limits (maximum 3 reply rounds) to prevent infinite communication loops.
Baselines
We compare C2C against two systematic baselines: (1) No Communication: agents work independently with task assignments but no communication, representing traditional parallel processing approaches; (2) Fixed Steps: agents communicate every 16 steps, simulating conventional project management practices. Both baselines use identical task decomposition and skill matching but lack C2C’s intelligent communication strategies.
Metrics
Following our evaluation framework, we report: task completion rate (percentage of tasks completed within time budget), average completion time (hours to finish successful tasks), communication cost (total time spent on communication activities), alignment score (average agent task alignment factor), and efficiency (ratio of productive work to total time investment). Each configuration is evaluated across all three task complexity tiers.
4.2 Main Results
As shown in Table˜1, which details the results for a 1M+4W team, C2C consistently demonstrates superior performance in key efficiency metrics, although all methods achieved a 100% task completion rate. The primary advantage of C2C is most evident in task completion time. On Complex tasks, C2C finishes in 24.75 hours, significantly faster than both No Communication (33.5 hours) and Fixed Steps (36.25 hours). This trend holds for Medium tasks, where C2C is also the fastest.
(a) Alignment evolution over rounds
(b) Impact of communication type
(c) Communication heatmap
Figure 3: Analysis of the Alignment Factor (AF) dynamics. (a) Average alignment improves over successive communication rounds across all task complexities. (b) High intent communications like MEETINGS and HELP requests yield the largest gains in alignment. (c) The communication heatmap reveals a manager centric coordination pattern, with the manager (M1) acting as the central hub.
This time saving translates to higher efficiency. For Complex tasks, C2C’s efficiency score is 1.62, substantially outperforming Fixed Steps (1.10) and No Communication (1.19). Similarly, C2C achieves the highest alignment score on Medium (0.68) and Complex (0.55) tasks, confirming that its communication strategy effectively enhances agent understanding. While communication costs are comparable across strategies, C2C utilizes its communication budget more effectively to achieve better outcomes.
Figure 3 illustrates the dynamics of the alignment factor mechanism. Alignment scores start low as agents have minimal task understanding, then improve through strategic communications. Help requests and meetings yield the strongest alignment improvements, while progress updates are modest. The communication heatmap reveals the manager centric communication pattern.
Configuration Completion Time Comm/Agent Comm Cost Speedup
1M + 4W 13h 3.1 2.75 1.54
1M + 8W 11.25h 2.1 3.78 1.78
1M + 16W 10.25h 2.6 5.12 1.95
1M+8W (2 tasks) 21h 2.3 4.64 1.35
Table 2: Scalability and multi-task analysis showing communication cost and speedup across team configurations. Multi-task performance demonstrates effective workload balancing and resource sharing capabilities. Speedup is calculated as the ratio of the completion time under the "No Communication" baseline to the completion time of each respective configuration.
4.3 Scalability and Multi-Task Analysis
Table 2 demonstrates C2C’s effectiveness across team sizes and workload scenarios. The framework exhibits sub-linear scaling in communication cost: while team size increases to 3.4 times (5 to 17 agents), communication cost only increases 86%. This favorable scaling behavior stems from C2C’s intelligent communication routing and the alignment factor mechanism, which prioritizes high value interactions over communications.
Speedup analysis reveals that performance steadily improves as the number of agents increases, with the largest configuration tested (1M+16W) achieving the highest speedup of $\mathbf{1.95\times}$. This highlights the effective task parallelization and specialized skill utilization in larger teams. However, the incremental gain in speedup lessens as the team grows, suggesting that simply adding more workers yields progressively smaller advantages.
The multi-task evaluation with 1M+8W reveals C2C’s ability to handle concurrent workloads effectively. When processing 2 tasks simultaneously, completion time increases from 11.25h to 21h (an 87% increase), significantly better than naive linear scaling. Analysis of communication patterns shows that C2C naturally evolves hub-and-spoke topologies with managers as primary coordinators, avoiding the quadratic communication complexity that plagues peer-to-peer approaches. In multi-task scenarios, agents exhibit sophisticated context switching behavior, maintaining separate alignment factors per task and prioritizing communications based on overall workflow optimization.
(a) Time vs. Complexity.
(b) Communication Freq. vs. Complexity.
(c) Message Type Distribution.
(d) Communication Method Analysis.
Figure 4: Effect of task complexity under C2C (8/24/40 hours) and team configuration (1M+4W, 1M+8W, 1M+16W). (a) Completion time increases with task complexity; larger worker pools shorten time with diminishing returns. (b) Messages per hour rise with both complexity and team size. (c) Help messages dominate across settings; clarification and meeting shares grow with complexity, and progress updates appear from medium upward. (d) Communication channels shift from chat (simple) toward email (complex), with a modest increase in meetings.
4.4 Message Type and Decomposition
Message Type Analysis
Table 3 details the impact of different message types on collaboration. MEETING_INVITE messages provide the highest alignment gains (+0.27), followed by HELP_REQUEST (+0.15), directly correlating with task success by resolving critical blockers. In contrast, PROGRESS_UPDATE messages maintain synchronization with a neutral (0) impact on alignment in this context. This analysis validates C2C’s strategy of prioritizing communications that address specific knowledge gaps.
Message Type Frequency Avg Response Step Impact on Alignment
HELP_REQUEST 9 3 +0.15
CLARIFICATION 0.3 2 +0.10
PROGRESS_UPDATE 0.3 5 0
MEETING_INVITE 1 7 +0.27
RESPONSE 8
Table 3: Analysis of message types in the 1M+8W configuration on Medium tasks, showing the frequency and impact of each communication type.
Task Decomposition Quality
Effective collaboration begins with high quality task decomposition. We compare C2C’s adaptive decomposition strategy against manual, naive LLM, and hierarchical methods in Table 4. C2C’s approach, which considers agent skills and workload during decomposition, achieves superior subtask clarity and higher AF. This leads to 14% increase in worker utilization rate compared to naive LLM decomposition, demonstrating that context aware planning is critical for multi-agent efficiency.
Decomposition Method Subtask Clarity Alignment Factor
Manual 1.00 0.70
LLM-naive 0.72 0.58
Hierarchical 0.89 0.64
C2C Adaptive 0.95 0.68
Table 4: Comparison of task decomposition quality. Subtask Clarity is rated by an LLM judge (1.0 is best). C2C’s adaptive method approaches manual quality.
4.5 Communication Pattern Analysis
Figure 4 shows how communication varies with task complexity and team size (1M+4W, 1M+8W, 1M+16W). (a) Completion time increases with task complexity; adding workers shortens time but with diminishing returns. (b) Message intensity (messages/hour) rises with complexity and team size.
Message Types
In (c), simple tasks are dominated by help and clarification (about 66% and 34%, respectively). At medium complexity, help remains the majority ($\approx$83%), while meetings appear ($\approx$8%). For complex tasks, the mix diversifies: help $\approx$70%, and meetings $\approx$13% become dominant. C2C allocates most messages to high value help requests and escalates to meetings only when the expected coordination benefit exceeds the cost.
Channels and Latency
Panel (d) indicates a shift from chat toward email as complexity grows, with a modest rise in meetings. The per channel response times for chat and email stay roughly flat across complexity levels (reduce slightly), whereas meeting latency increases for complex tasks.
These patterns align with the logic of C2C: as tasks become more complex, agents continue to seek help most of the time but increasingly use meetings, while clarification needs drop. The engine selects higher yield (though costlier) channels when needed, while larger teams reduce completion time without eliminating the upward pressure on communication volume with complexity.
5 Conclusion
In this paper, we present Communication to Completion, a multi-agent framework that treats communication as an optimizable resource and quantifies task understanding through the Alignment Factor within a Sequential Action Framework. By treating alignment as a quantifiable and dynamic variable, C2C enables agents to autonomously determine when and how to communicate.
Our experiments demonstrate that this alignment-driven approach adapts naturally to task complexity, with agents communicating more frequently and through richer channels when facing difficult problems while minimizing overhead on simpler tasks. Across diverse task complexities and team sizes, C2C consistently achieves high task completion while reducing completion time and improving efficiency relative to baselines. The framework produces interpretable patterns, including alignment factor trajectories, communication type distributions, and collaboration network structures, which bridge agent behavior and coordination theory to provide design insights for multi-agent systems.
Limitations
While the C2C framework demonstrates significant performance gains, we acknowledge several limitations that offer avenues for future research. The findings are derived from a controlled simulation, and the framework’s performance in unpredictable, real-world environments is yet to be validated. The experiments were confined to the software engineering domain, so the C2C’s effectiveness may not generalize to other collaborative fields. Furthermore, the Alignment Factor (AF) relies on an LLM’s evaluation, which introduces potential subjectivity and bias.
Appendix A Tasks Details
This part provide example task prompts used in our simulations. Each prompt specifies the description, time budget, and required skills.
Simple Task description: "Fix five independent bugs across modules: login validation, data parsing, UI rendering glitches, API timeout handling, and a database connection leak. No cross-dependencies." hours: 8.0 skills: ["backend", "frontend", "database", "api", "testing"]
Medium Task description: "Integrate an external API with authentication, including token management and error/latency handling; deliver minimal usage docs." hours: 24.0 skills: ["backend", "api", "authentication", "oauth", "testing", "documentation"]
Hard Task description: "Build a user authentication service covering registration, login, password reset, OAuth 2.0 sign-in, JWT issuance/refresh, session management, and security hardening." hours: 40.0 skills: ["backend", "security", "database", "oauth", "authentication", "frontend", "testing"]
Appendix B Implementation Details
We provide prompts used in our experiments in this section.
Decompose a task into subtasks Decompose this task into subtasks:
Task: {task.description}
Estimated hours: {task.estimated_hours}
Required skills: {’, ’.join(task.required_skills)}
Team members:
{self._format_team(team)}
{manager_note}
Create {teamsize} subtasks that:
1. Can be worked on independently or with minimal dependencies
2. Match team members’ skills
3. Are reasonably sized (total hours should be close to the original task’s estimated hours)
4. Cover all aspects of the original task
Return JSON: {{ "subtasks": [ {{ "description": "Clear description of what needs to be done", "estimated_hours": number, "required_skills": ["skill1", "skill2"], "suggested_assignee": "team member_name (can be any team member including Manager)", "dependencies": [] // indices of subtasks this depends on }} ], "decomposition_rationale": "Brief explanation of your decomposition strategy" }}
Figure 5: Prompt for the manager to decompose a give task.
Intention generation You are {context.agent.name}, a {context.agent.role.value}. Current situation: Tasks: {self._format_tasks_for_prompt(context.current_tasks[:3], context)} Note: Alignment affects work efficiency.(max: 1.0) {message_info} {last_action_info} Analyze your situation and choose your primary intention for this step: 1. CONTINUE_TASK - Continue working on current tasks 2. CHECK_MESSAGES - Check and potentially respond to messages 3. REQUEST_HELP - Ask for other agents help 4. NEED_CLARIFICATION - Need clarification on task requirements 5. REPORT_PROGRESS - Report progress to manager 6. SCHEDULE_MEETING - Schedule a meeting IMPORTANT Decision Factors: - If has MEETING_START: Almost always CHECK_MESSAGES (meeting is starting!) - If has MEETING_INVITE: Strongly consider CHECK_MESSAGES (need to RSVP) - If stuck on task for long: Consider REQUEST_HELP or NEED_CLARIFICATION - Balance responsiveness with productivity Return JSON: {{"intention": "INTENTION_NAME", "reasoning": "explanation"}}
Figure 6: Prompt agents use to generate their intentions at each step.
Decompose a task into subtasks You are evaluating how much a received message helps an worker understand their task better. Task Information: - Task ID: {task_id} - Description: {task.description} - Required Skills: {’, ’.join(task.required_skills) if task.required_skills else Not specified’} - Current Progress: {task.actual_hours:.1f}/{task.estimated_hours:.1f} hours - Current Alignment Factor: {current_alignment:.2f} ({current_alignment*100:.0f}% efficiency) Message Type: {message.message_type.value if hasattr(message, message_type’) else Unknown’} From Agent: {message.from_agent_id} {’Original Request (sent by me):’ if original_request else Context:’} {original_request.content if original_request else This is a proactive message or the original request is not available.’} Reply Received: {message.content} Alignment Factor Guidelines: - Range: 0.01 (1% efficiency) to 1.0 (100% efficiency) - Current value: {current_alignment:.2f} - Alignment can increase OR decrease based on communication quality - Clear and helpful communication should improve understanding - Confusing, contradictory, or misleading information may reduce understanding - Consider the overall impact on task clarity and execution confidence Consider: 1. How directly the reply addresses the original request/confusion 2. How actionable and specific the information is 3. Whether critical blockers were resolved 4. The completeness of the response 5. Diminishing returns (harder to improve as alignment approaches 1.0) Return a JSON response: {{ "new_alignment_factor": <float between 0.01 and 1.0>, "change": <float, positive for increase, negative for decrease>, "reasoning": "<brief explanation of why this change in understanding>" }}
Figure 7: Prompt for updating alignment factor.

### v2

@@
 Abstract
-Teamwork in workspace for complex tasks requires diverse communication strategies, but current multi-agent LLM systems lack systematic frameworks for task oriented communication.
-We introduce Communication to Completion (C2C), a scalable framework that addresses this gap through two key innovations: (1) the Alignment Factor (AF), a novel metric quantifying agent task alignment that directly impacts work efficiency, and (2) a Sequential Action Framework that integrates stepwise execution with intelligent communication decisions.
-C2C enables agents to make cost aware communication choices, dynamically improving task understanding through targeted interactions.
-We evaluated C2C on realistic coding workflows across three complexity tiers and team sizes from 5 to 17 agents, comparing against no communication and fixed steps baselines.
-The results show that C2C reduces the task completion time by about 40% with acceptable communication costs.
-The framework completes all tasks successfully in standard configurations and maintains effectiveness at scale.
-C2C establishes both a theoretical foundation for measuring communication effectiveness in multi-agent systems and a practical framework for complex collaborative tasks.
+Multi-agent LLM systems have demonstrated impressive capabilities in complex collaborative tasks, yet most frameworks treat communication as instantaneous and free, overlooking a fundamental constraint in real world teamwork, collaboration cost.
+We propose a scalable framework implemented via Communication to Completion (C2C), which explicitly models communication as a constrained resource with realistic temporal costs.
+We introduce the Alignment Factor (AF), a dynamic metric inspired by Shared Mental Models, to quantify the link between task understanding and work efficiency.
+Through experiments on 15 software engineering workflows spanning three complexity tiers and team sizes from 5 to 17 agents, we demonstrate that cost-aware strategies achieve over 40% higher efficiency compared to unconstrained interaction.
+Our analysis reveals emergent coordination patterns: agents naturally adopt manager centric hub-and-spoke topologies, strategically escalate from asynchronous to synchronous channels based on complexity, and prioritize high value help requests.
+These patterns remain consistent across multiple frontier models (GPT-5.2, Claude Sonnet 4.5, Gemini 2.5 Pro).
+This study moves beyond simple agent construction, offering a theoretical foundation for quantifying and optimizing the dynamics of collaboration in future digital workplaces.
 Communication to Completion: Modeling Collaborative Workflows with Intelligent Multi-Agent Communication
@@
 1 Introduction
-Figure 1: Communication to Completion architecture: a five-layer stack simulation engine
-Modern organizations run on teams: people plan, divide, and verify work under deadlines while juggling synchronous meetings, asynchronous chats/emails, and limited attention.
-Extending this paradigm to AI, coordinating teams of LLM agents to execute complex, long-horizon tasks (e.g., software development) remains a grand challenge (Park et al., 2023; Hong et al., 2024; Qian et al., 2023).
-Effective collaboration hinges on communication: too much creates coordination overhead; too little yields misalignment and rework.
-Recent multi-agent frameworks demonstrate that structured dialogues and role specialization can transform general purpose LLMs into cooperative problem solvers across software engineering, data analysis, and decision support (Wu et al., 2024; Hong et al., 2024; Qian et al., 2023; Li et al., 2023).
-Beyond mere tool use, conversation itself decomposes tasks, aligns partial knowledge, and arbitrates proposals, often outperforming single agent pipelines on complex problems (de Souza Santos and Ralph, 2022; Khan et al., 2024; Yao et al., 2023a; Besta et al., 2024).
-Yet many systems still schedule or trigger communication via fixed or purely reactive heuristics that ignore the evolving trade-off between coordination cost and task progress (Wu et al., 2024; Chen et al., 2023).
-Meanwhile, empirical studies of software teams warn that coordination costs (especially meetings), can dominate delivery time in modern distributed work, with a post-2020 shift toward heavier synchronous load and changing patterns across chat, email, and meetings (Mok et al., 2023; Yang et al., 2022).
-These observations motivate treating communication as a first class resource to be scheduled, routed, and evaluated rather than as a byproduct of agent reasoning, thus leading to a central open question: how can agents learn to communicate strategically to maximize collaborative benefit while minimizing coordination costs?
-In this work, we address the above question by casting multi-agent collaboration as cost communication and optimizing it accordingly. Our contributions are:
-  • We introduce Communication to Completion (C2C).
-A scalable framework models when, whom, and how to communicate as part of execution and instantiate it with a Sequential Action Framework (SAF) that synchronizes agent actions into discrete timesteps for deterministic, reproducible collaboration.
-  • We define Alignment Factor (AF), a metric that quantifies each agent’s task understanding and directly modulates work efficiency.
-Communications update AF, and AF in turn scales work rate, linking conversation to progress and yielding interpretable traces.
-  • We demonstrate through comprehensive experiments spanning complexities and team sizes, C2C lowers completion time and improves work efficiency compared to baselines.
-2 Related Work
-Multi-Agent Collaboration Frameworks
-LLM agents are increasingly organized as teams rather than monoliths.
-Social simulation work demonstrates emergent, human-like collective behavior (e.g., the “Smallville” agents of Park et al.
-(2023) and the recent large scale AgentSociety simulation (Piao et al., 2025)).
-For goal oriented collaboration—particularly software development—frameworks such as MetaGPT (Hong et al., 2024) and ChatDev (Qian et al., 2023) impose predefined roles and SOPs that mirror corporate workflows.
-AutoGen (Wu et al., 2024) offers a flexible substrate for composing conversable agents, while AgentVerse (Chen et al., 2023) and OpenAgents (Xie et al., 2023) provide extensible environments for building and deploying general purpose agents.
-Outside pure research prototypes, platforms like OpenHands operationalize multi/single agent development loops in realistic toolchains (Wang et al., 2025).
-In parallel, role-playing and deliberation structures further organize collaboration within conversations (Li et al., 2023; Yao et al., 2023a; Besta et al., 2024), and new evaluations emphasize interactive, multi-environment testing and repository-level search and planning (Liu et al., 2023a; Antoniades et al., 2024).
-Our work differs in elevating communication from a procedural step to an optimizable resource, explicitly linking when/what to communicate with collaborative progress via an alignment driven mechanism.
-Communication in Multi-Agent Systems
-Communication has long been central in multi-agent learning.
-Classic MARL studies learned protocols end-to-end under partial observability (Foerster et al., 2016; Sukhbaatar et al., 2016).
-Recent advances focus on efficient or sparse communication to reduce overhead, e.g., decentralized scheduling and model-based message estimation (Han et al., 2023; Liu et al., 2023c; Kim et al., 2019; Singh et al., 2018).
-In LLM agent settings, natural-language messaging carries high-level, semantic content, and methods increasingly structure agent talk through reasoning–acting loops (ReAct) (Yao et al., 2023b), self improvement with textual feedback (Reflexion) (Shinn et al., 2023), and iterative self-refinement (Madaan et al., 2023).
-Multi-agent debate further uses conversational arbitration to improve correctness and planning (Du et al., 2023).
-Yet many systems still trigger communication on fixed or purely reactive heuristics.
-We instead propose an alignment driven policy that decides when, whom and how to communicate based on a measurable, task grounded alignment signal.
+Figure 1: Overview of the C2C simulation framework.
+Agents autonomously select between execution and communication channels by weighing the Alignment Factor (AF) against modeled latencies, forming a closed-loop system where communication is treated as a constrained resource.
+Effective teamwork in software development hinges on strategic communication.
+Teams must continuously balance the value of information sharing against its costs: synchronous meetings consume collective time, asynchronous messages introduce delays, and frequent interruptions fragment attention (Mok et al., 2023; Yang et al., 2022).
+Understanding when, with whom and how to communicate remains a central challenge in organizational research (Marlow et al., 2018).
+However, systematically studying these dynamics in real teams is notoriously difficult due to the inability to control variables or replay scenarios.
+Recent advances in large language models (LLMs) offer a new approach to simulate workplace behaviors.
+Just as Park et al.
+(2023) demonstrated that LLM agents can exhibit believable human-like social behavior in simulated communities, we propose using multi-agent LLM systems as a controlled simulation environment for investigating workplace collaboration patterns.
+Despite this potential, current multi-agent frameworks are not well-suited for realistic workplace simulation.
+Systems such as AutoGen and MetaGPT typically abstract away communication costs, effectively assuming instantaneous information exchange (Wu et al., 2024; Hong et al., 2024).
+As a result, agents can exchange unlimited tokens without explicitly modeling temporal latency or cognitive costs (e.g., context switching) inherent in real workflows.
+Consequently, these frameworks cannot capture the fundamental trade-off between alignment and coordination cost, and may encourage excessive communication that would be unsustainable for human teams.
+To bridge this gap, we introduce Communication to Completion (C2C), a framework designed to simulate collaborative workflows under realistic constraints.
+C2C incorporates a discrete event simulation engine that explicitly models communication costs, enabling communication to be treated as a scarce resource to be optimized.
+Within this environment, we propose the Alignment Factor (AF), a dynamic metric inspired by the concept of Shared Mental Models (Cannon-Bowers et al., 1993), to quantify how effectively agents build and maintain shared task understanding over time.
+We use AF both as an evaluation metric and as a control signal to guide when additional communication is worth its cost.
+Our contributions are as follows:
+  • We establish a computational sandbox for multi-agent collaboration that moves beyond instantaneous communication assumptions, enabling the study of agent behaviors under realistic temporal and cognitive constraints.
+  • We introduce the Alignment Factor as a quantifiable metric for shared task understanding.
+We show that optimizing for AF enables agents to dynamically balance the benefits of information sharing against the costs of coordination.
+  • Through experiments on complex software engineering tasks, we demonstrate that cost aware communication strategies outperform unconstrained interaction in efficiency.
+We further analyze emergent coordination patterns, showing how constraints induce more structured communication topologies.
 Figure 2: C2C timeline on a software bug-fix task.
@@
 Vertical dashed dividers denote discrete timesteps in the Sequential Action Framework; the cycle repeats until completion (Step N).
-Quantifying Coordination and Alignment
-Coordination costs are a foundational concern in collaborative systems (Malone, 1988; Brooks Jr, 1995).
-However, evaluations of LLM multi agents often emphasize final outcomes (e.g., pass rates on benchmarks) over process efficiency.
-In software engineering, SWE-bench exposes the gap between realistic issue resolution and model capabilities (Jimenez et al., 2023), while SWE-agent (Yang et al., 2024) and OpenHands (Wang et al., 2025) reveal the importance of tool use and orchestration—but still provide limited real-time signals about team alignment.
-Complementary efforts argue for richer process traces and interactive environments to analyze coordination choices (Liu et al., 2023a; Antoniades et al., 2024), and work on LLM-as-judge warns that automated evaluators may introduce bias at the process level (Liu et al., 2023b; Chen et al., 2024; Tan et al., 2024).
-We contribute the Alignment Factor (AF) as a dynamic, real-time proxy for shared task understanding that both modulates work efficiency and directly informs agents’ communication decisions, thereby connecting conversational actions to tangible task outcomes.
+2 Related Work
+2.1 Multi-Agent Collaboration Frameworks
+Recent research on LLM multi-agent systems broadly spans two lines of work: task oriented execution frameworks and behavioral simulations.
+Task oriented frameworks such as MetaGPT (Hong et al., 2024), ChatDev (Qian et al., 2023), and AutoGen (Wu et al., 2024) employ role structures, Standard Operating Procedures (SOPs), or conversable agents to solve complex engineering tasks.
+While effective for end-to-end task completion, these systems typically do not explicitly model communication as a constrained resource: interaction is commonly treated as low friction message passing, without representing temporal latency or cognitive overhead that shape real world teamwork.
+Conversely, behavioral simulations such as Generative Agents (Park et al., 2023) and AgentSociety (Piao et al., 2025) focus on reproducing believable human social dynamics and daily routines, but are less tied to goal directed engineering workflows.
+Our work bridges these directions by adopting the simulation perspective while targeting task-oriented collaboration: we explicitly model coordination constraints and costs to enable controlled study of communication strategies under resource limitations.
+2.2 Communication Costs and Constraints
+Communication overhead is a foundational concern in organizational theory and collaborative systems, and has also been widely studied in multi-agent reinforcement learning (MARL).
+In MARL, agents learn to selectively communicate or gate messages to reduce bandwidth usage and coordination overhead (Singh et al., 2018; Kim et al., 2019).
+In LLM-based multi-agent systems, communication is predominantly natural language and is often treated as part of the reasoning process rather than an explicit resource.
+Although recent work structures agent interaction through protocols such as debate and other deliberative exchanges (Du et al., 2023; Khan et al., 2024), mainstream frameworks still rarely represent costs such as asynchronous delay or context switch overhead.
+Empirical studies in software engineering further suggest that coordination costs (e.g., synchronous meetings and scheduling friction) can dominate delivery time in distributed teams (Mok et al., 2023; Yang et al., 2022).
+Our framework explicitly models such costs via a discrete event simulation engine, distinguishing our approach from systems that view communication primarily as a byproduct of reasoning.
+2.3 Shared Mental Models and Alignment
+Effective teamwork relies on Shared Mental Models (SMMs), the overlapping knowledge structures held by team members that support coordinated action (Cannon-Bowers et al., 1993; Mathieu et al., 2000).
+In human teams, misalignment can lead to “process loss”, where effort is diverted to resolving misunderstandings rather than executing tasks.
+For LLM agents, assessing alignment remains challenging.
+Most evaluations emphasize outcome correctness (e.g., pass@1 on SWE-bench (Jimenez et al., 2023; Yang et al., 2024)) rather than process quality.
+While recent benchmarks and analyses examine interaction trajectories (Liu et al., 2023a; Yang et al., 2024), they do not provide a dynamic measure of consensus during collaboration.
+We introduce the Alignment Factor (AF) as a computational proxy for SMM, enabling quantitative measurement of how communication updates shared context and how evolving alignment relates to downstream execution efficiency.
 3 C2C Framework
-Communication to Completion is a time-stepped framework for deciding when, whom, and how to communicate in multi-agent collaboration.
-Figure 2 provides an overview of its mechanism using a software task.
-We organize the discussion as follows: §3.1 introduces the Sequential Action Framework that ensures deterministic execution, §3.2 describes the Alignment Factor mechanism for quantifying task understanding, §3.3 details how agents make communication decisions, §3.4 explains hierarchical task decomposition and tracking, and §3.5 presents the intention-based agent decision making that integrates these components.
-3.1 Sequential Action Framework
+Communication to Completion is a simulation framework for deciding when, with whom, and how to communicate in multi-agent collaboration.
+Figure 2 provides an overview of the mechanism.
+We organize the discussion as follows: §3.1 introduces the discrete event engine that ensures deterministic execution, §3.2 describes the Alignment Factor for quantifying task understanding, §3.3 details how agents make cost-aware communication decisions, and §3.4 explains hierarchical task decomposition.
+3.1 Discrete Event Simulation Engine
 Concurrent multi-agent execution often introduces temporal ambiguity.
-The Sequential Action Framework (SAF) constrains each agent to exactly one action per timestep, yielding deterministic state transitions and reproducible analysis.
+To ensure reproducibility, we employ a discrete event simulation engine that constrains each agent to exactly one action per timestep, yielding deterministic state transitions.
 Actions
-SAF defines four actions capturing common collaborative behavior that capture common collaborative behavior: work (task execution with time cost), communicate (compose/send a message), reply (respond to an incoming message), and meeting (synchronous group discussion).
+The engine defines four actions capturing common collaborative behavior: work (task execution with time cost), communicate (compose/send a message), reply (respond to an incoming message), and meeting (synchronous group discussion).
 Each action is temporally bounded and commits a single transition.
@@
 $\forall i,j:\;a_{i}^{t}\;\text{completes before}\;a_{j}^{t+1}\;\text{starts.}$ (1)
-Forward-only Delivery.
+Forward-only Delivery
 Messages sent at $t$ are delivered at t+1t{+}1, preventing instantaneous feedback loops and enforcing causal consistency.
@@
 3.2 Alignment Factor
-The Alignment Factor (AF) measures an agent’s task-specific understanding and directly modulates effective productivity.
+The Alignment Factor (AF) serves as a proxy for the Shared Mental Model (Cannon-Bowers et al., 1993), measuring an agent’s task specific understanding.
 Unlike static capability scores, AF evolves through communication.
 For agent $i$ on task $j$, $AF_{i,j}\in[0.01,1.00]$ denotes comprehension quality.
-Values near $0.01$ indicate limited understanding; values near $1.00$ indicate mastery.
-We initialize $AF_{i,j}{=}0.3$ to reflect the initial understanding.
-When agent $i$ receives reply about task $j$ from agent $k$, AF is updated based on the communication’s impact on task understanding.
-An LLM evaluates the message content, agent $i$’s current task context, and the relevance of the received information to determine the new alignment:
+We initialize $AF_{i,j}{=}0.3$ to reflect the initial ambiguity.
+When agent $i$ receives information, the AF is updated based on the communication’s quality.
+To rigorously measure this without agent bias, the update is computed by the simulation environment (using ground-truth context) rather than the agent itself:
 $AF_{i,j}^{\text{new}}=\min\!\bigl(1.0,\;AF_{i,j}^{\text{old}}+\Delta_{\text{eval}}\bigr),$ (2)
-where $\Delta_{\text{eval}}\in[0,0.5]$ is computed by the LLM based on: (i) how well the message addresses the agent’s knowledge gaps, (ii) the relevance of the information to the specific task requirements, and (iii) the clarity and actionability of the guidance provided.
-Different message types naturally yield different impacts, help requests addressing critical blockers typically produce larger alignment gains.
+where $\Delta_{\text{eval}}\in[0,0.5]$ is computed based on: (i) knowledge gap resolution, (ii) relevance to task requirements, and (iii) actionability.
 If agent $i$ spends $h$ hours on task $j$, the effective progress is
 $\text{EffectiveProgress}\;=\;h\cdot AF_{i,j},$ (3)
-A low AF impedes an agent’s progress, creating a natural incentive for it to seek clarification and coordination before investing substantial effort.
+A low AF impedes progress, creating a natural incentive to seek clarification despite the communication cost.
 3.3 Agent Communication Decisions
-Agents autonomously decide when, with whom, and how to communicate based on their current task state and situational awareness, without relying on fixed thresholds or rigid scheduling.
-For clarity, we decompose this decision process into four key components: initiating the communication, selecting the recipients, choosing the appropriate channel, and composing the message content.
+Agents autonomously decide communication strategies based on situational awareness.
+Crucially, this decision process is cost aware, balancing the need for alignment against simulated channel constraints.
 Communication Initiation
-Agents may initiate communication in response to various task situations: encountering technical difficulties (help_request), facing unclear requirements (need_clarification), identifying coordination needs with team members (meeting_invite), or reaching milestones that warrant status updates (progress_update).
-The decision to communicate emerges from the agent’s assessment of its current alignment and task complexity.
+Agents may initiate communication in response to task situations such as technical difficulties (help_request) or uncoordinated dependencies (meeting_invite).
+The decision emerges from the agent’s assessment of its current AF versus the time cost of interaction.
 Recipient Selection
-When initiating communication, agents consider multiple factors in selecting recipients: skill relevance to the issue, existing task dependencies, and historical interaction.
-The selection process reflects realistic collaboration patterns observed in human teams.
-Channel Selection
-Agents choose among three communication channels.
-Chat suits quick questions and simple clarifications with rapid response times.
-Email handles detailed explanations with longer response windows.
-Meetings address complex coordination needs requiring synchronous discussion among multiple participants.
+Agents select recipients based on skill relevance and historical interaction, reflecting realistic collaboration patterns.
+Channel Selection and Costs
+Agents operate on a utility maximization principle. We model the utility $U_{i,ch}$ for agent $i$ choosing channel $ch$ as:
+$U_{i,ch}=w_{1}\cdot\mathcal{S}(D_{task},B_{ch})-w_{2}\cdot\mathcal{C}(T_{lat},N)+\rho_{i,ch}$ (4)
+where:
+  • Task Suitability ($\mathcal{S}$): Measures how well the channel’s bandwidth $B_{ch}$ matches the current task complexity $D_{task}$.
+Complex tasks yield higher utility in high bandwidth channels (Meetings), whereas simple tasks (e.g., status check) favor low ones.
+  • Time Cost ($\mathcal{C}$): Penalizes transmission latency $T_{lat}$ and synchronization overhead for $N$ participants.
+  • Agent Preference ($\rho_{i,ch}$): Captures agent specific biases, $\rho>0$ for preferred channels.
+This composite formulation ensures that agents do not merely seek the fastest route, but dynamically select the most appropriate channel given the task’s nature and their role constraints.
 Message Composition
-Communication content is generated based on the agent’s current context, including task description, encountered difficulties, current alignment factor, and relevant technical details.
-Messages are composed to be informative and actionable, providing recipients with sufficient context to offer effective assistance.
+Communication content is generated based on the agent’s current context, including task description and alignment gaps. Messages are composed to be informative.
 3.4 Hierarchical Task Management
@@
 At each timestep, the agent constructs a comprehensive context that includes: (i) currently assigned tasks and their completion status, (ii) alignment factors for each assignment indicating task understanding, (iii) recent communications including pending requests and received guidance, and (iv) team state including teammate skills and availability
+Metric Complexity No Comm Fixed Steps Free Comm C2C (Ours)
+Task Completion Rate (%) Simple $100$ $100$ $96\pm 8$ $100$
+Medium $100$ $100$ $100$ $100$
+Complex $100$ $100$ $100$ $100$
+Avg Completion Time (hours) $\downarrow$ Simple $7.00$ $\mathbf{5.00\pm 0.3}$ $6.00\pm 0.4$ $5.50\pm 0.25$
+Medium $20.0$ $14.75\pm 0.75$ $14.5\pm 1.1$ $\mathbf{13.00\pm 0.9}$
+Complex $33.5$ $36.25\pm 3.25$ $28.0\pm 2.8$ $\mathbf{24.75\pm 1.5}$
+Communication Cost (hours) $\downarrow$ Simple $2.03\pm 0.1$ $3.20\pm 0.6$ $\mathbf{1.94\pm 0.2}$
+Medium $2.75\pm 0.2$ $5.40\pm 1.5$ $\mathbf{3.26\pm 0.2}$
+Complex $8.12\pm 0.9$ $14.55\pm 3.2$ $\mathbf{7.02\pm 0.8}$
+Alignment Score (AF) $\uparrow$ Simple $0.30$ $0.55\pm 0.05$ $\mathbf{0.60\pm 0.04}$ $0.51\pm 0.03$
+Medium $0.30$ $0.59\pm 0.06$ $0.63\pm 0.05$ $\mathbf{0.68\pm 0.04}$
+Complex $0.30$ $0.53\pm 0.08$ $\mathbf{0.65\pm 0.07}$ $0.55\pm 0.05$
+Efficiency (Work/Time) $\uparrow$ Simple $1.14$ $\mathbf{1.60\pm 0.1}$ $1.33\pm 0.7$ $1.45\pm 0.07$
+Medium $1.20$ $1.63\pm 0.1$ $1.66\pm 0.12$ $\mathbf{1.85\pm 0.13}$
+Complex $1.19$ $1.10\pm 0.09$ $1.43\pm 0.14$ $\mathbf{1.62\pm 0.1}$
+Table 1: Main results comparing collaboration policies on GPT-4o agents.
+Free Comm represents an unconstrained baseline where agents communicate without simulated delays, resulting in faster raw completion but high communication costs.
+Values are reported as mean $\pm$ standard deviation over 5 runs.
+C2C achieves the highest Efficiency (task progress normalized by total time investment) on Medium and Complex tasks, demonstrating the value of cost-aware strategy.
+($\downarrow$: lower is better, $\uparrow$: higher is better).
 Intention
-Given this context, the agent uses an LLM to carefully reason about the current situation and generate an intention, a high-level decision about what action to take next.
-The generated intention is then translated into a concrete action within the SAF framework:
-  • Work: Continue task execution with effectiveness modulated by current AF.
-  • Help: Compose the message identifying difficulties and required expertise (HELP_REQUEST).
-  • Clarification: Request additional task details or requirement explanations (NEED_CLARIFICATION).
-  • Coordination: Propose synchronous discussion for complex collaboration needs (MEETING_INVITE).
-  • Reporting: Share subtasks progress with manager (PROGRESS_REPORT).
+Given this context, the agent uses an LLM to generate an intention. The generated intention translates into actions within the simulation:
+  • Work: Continue task execution (efficiency modulated by AF).
+  • Help/Clarification: Request assistance (incurs communication cost).
+  • Coordination: Propose meetings (high synchronization cost).
+  • Reporting: Share progress.
 When generating communication intentions, the agent also determines appropriate recipients and selects suitable channels (chat, email, meeting) based on urgency and complexity.
@@
 4.1 Experimental Setup
-Metric Complexity No Communication Fixed Steps C2C
-Task Completion Rate (%) Simple 100 100 100
-Medium 100 100 100
-Complex 100 100 100
-Avg Completion Time (hours) Simple 7 5 5.5
-Medium 20 14.75 13
-Complex 33.5 36.25 24.75
-Communication Cost Simple 2.03 1.94
-Medium 2.75 3.26
-Complex 8.12 7.02
-Alignment Score (AF) Simple 0.3 0.55 0.51
-Medium 0.3 0.59 0.68
-Complex 0.3 0.53 0.55
-Efficiency Simple 1.14 1.6 1.45
-Medium 1.20 1.63 1.85
-Complex 1.19 1.10 1.62
-Table 1: Comparison of three policies for a 1M+4W team across task complexities (Simple=8h, Medium=24h, Complex=40h).
-We report Task Completion Rate (%), Avg Completion Time (hours), Communication Cost (total communication time), Alignment Score (mean AF over the run; higher indicates better shared understanding), and Efficiency == completed work / total time.
-C2C achieves the best time and efficiency on Medium and Complex with comparable communication cost.
-Bold marks the best value within each row; "-" denotes not applicable.
-To evaluate the performance of C2C on realistic tasks, we implement experiments on software engineering workflows across three complexity tiers: Simple (8 hours, basic SWE operations), Medium (24 hours, API integration with authentication), and Complex (40 hours, full stack user authentication system); see Appendix A for details.
-Each task requires diverse skills including front-end development, back-end systems, database management, security implementation, and testing.
-Tasks are decomposed into several subtasks according to task complexity and team size with explicit skill requirements and dependencies, mirroring real-world software development scenarios.
-Team Configurations
-We systematically vary team sizes from 5 to 17 agents: 1M+4W (5 agents), 1M+8W (9 agents), and 1M+16W (17 agents), where M denotes manager and W denotes worker.
-Additionally, we evaluate multi-task scenarios with 1M+8W handling 2 concurrent tasks to assess workload distribution and task interference effects.
-Each agent is assigned 2–4 complementary skills from a skill pool, including skills like frontend, backend, and database.
-Models and Implementation
-All agents are powered by GPT-4o with temperature 0.7 for human like decision making.
-The Sequential Action Framework operates with 0.25 hour time steps over a maximum of 160 simulation steps (40 hours).
-Communication costs are calculated using realistic time models: email drafting (9 minutes base + content length), chat messages (3 minutes base), meetings (30 minutes minimum + preparation time + number of participants).
-We implement thread-depth limits (maximum 3 reply rounds) to prevent infinite communication loops.
-Baselines
-We compare C2C against two systematic baselines: (1) No Communication: agents work independently with task assignments but no communication, representing traditional parallel processing approaches; (2) Fixed Steps: agents communicate every 16 steps, simulating conventional project management practices.
-Both baselines use identical task decomposition and skill matching but lack C2C’s intelligent communication strategies.
-Metrics
-Following our evaluation framework, we report: task completion rate (percentage of tasks completed within time budget), average completion time (hours to finish successful tasks), communication cost (total time spent on communication activities), alignment score (average agent task alignment factor), and efficiency (ratio of productive work to total time investment).
-Each configuration is evaluated across all three task complexity tiers.
-4.2 Main Results
-As shown in Table˜1, which details the results for a 1M+4W team, C2C consistently demonstrates superior performance in key efficiency metrics, although all methods achieved a 100% task completion rate.
-The primary advantage of C2C is most evident in task completion time.
-On Complex tasks, C2C finishes in 24.75 hours, significantly faster than both No Communication (33.5 hours) and Fixed Steps (36.25 hours).
-This trend holds for Medium tasks, where C2C is also the fastest.
+To evaluate the performance of C2C on realistic tasks, we constructed a benchmark dataset consisting of 15 software engineering workflows across three complexity tiers (5 tasks per tier) to ensure robustness across different business domains (see Appendix A for full task list).
+For clarity in our detailed ablation studies, we report metrics from the representative task of each tier:
+  • Simple (8 hours): Basic SWE operations (e.g., bug fixes, refactoring).
+  • Medium (24 hours): Module integration (e.g., API integration with authentication).
+  • Complex (40 hours): Full-stack system design (e.g., user authentication service).
+Each task requires diverse skills including front-end development, back-end systems, database management, and testing.
+Tasks are decomposed into subtasks with explicit dependencies, mirroring real-world software development scenarios.
 (a) Alignment evolution over rounds
@@
 (c) The communication heatmap reveals a manager centric coordination pattern, with the manager (M1) acting as the central hub.
-This time saving translates to higher efficiency.
-For Complex tasks, C2C’s efficiency score is 1.62, substantially outperforming Fixed Steps (1.10) and No Communication (1.19).
-Similarly, C2C achieves the highest alignment score on Medium (0.68) and Complex (0.55) tasks, confirming that its communication strategy effectively enhances agent understanding.
-While communication costs are comparable across strategies, C2C utilizes its communication budget more effectively to achieve better outcomes.
-Figure 3 illustrates the dynamics of the alignment factor mechanism.
-Alignment scores start low as agents have minimal task understanding, then improve through strategic communications.
-Help requests and meetings yield the strongest alignment improvements, while progress updates are modest.
-The communication heatmap reveals the manager centric communication pattern.
+Team Configurations
+We systematically vary team sizes from 5 to 17 agents: 1M+4W (5 agents), 1M+8W (9 agents), and 1M+16W (17 agents), where M denotes manager and W denotes worker.
+Additionally, we evaluate multi-task scenarios with 1M+8W handling 2 concurrent tasks.
+Models and Implementation
+Agents are powered by GPT-4o (temperature 0.7).
+We also evaluate Claude 3.5 Sonnet to test generalization.
+The discrete event simulation engine operates with 0.25-hour time steps over a maximum of 160 steps.
+Communication costs model real world constraints: Email (9 min delay + drafting), Chat (3 min), and Meetings (30 min blocking cost).
+Baselines
+We compare C2C against three baselines: (1) No Communication: Agents work independently.
+(2) Fixed Steps: Agents communicate on a rigid schedule (every 16 steps).
+(3) Free Communication: An unconstrained baseline where agents communicate without simulated latency, representing idealized interaction.
+Metrics
+We report: task completion rate (percentage of tasks completed within time budget), average completion time (hours to finish successful tasks), communication cost (total time spent on communication activities), alignment score (average agent task alignment factor), and efficiency (ratio of productive work to total time investment).
+Each configuration is evaluated across all three task complexity tiers.
+4.2 Main Results
+Table 1 presents the performance comparison.
+While all methods achieve high completion rates, C2C consistently demonstrates superior Efficiency on Medium (1.85) and Complex (1.62) tasks.
+Comparison with the Free Communication baseline reveals the impact of communication overhead.
+Although Free Comm allows agents to communicate instantly, this leads to excessive interaction (e.g., 14.55 hours cost on Complex tasks vs.
+7.02 hours for C2C).
+This chatter distracts agents from execution, effectively slowing down the overall progress.
+C2C, by weighing the cost of communication against alignment gains, finds an optimal balance, engaging in just enough coordination to maintain high Alignment Factor (AF) without incurring high costs.
+To understand the dynamics behind this efficiency, Figure 3 analyzes the evolution of the Alignment Factor.
+As shown in Figure 3(a), alignment scores start low and improve progressively.
+Figure 3(b) highlights that high cost interactions like Meetings yield the largest alignment gains justifying their use for critical synchronization, whereas low cost messages provide smaller increments.
+The heatmap in Figure 3(c) confirms a manager coordination pattern, where the manager acts as a hub to distribute context efficiently.
+4.3 Generalization across Models
+To verify that our findings are not specific to GPT-4o, we evaluated C2C using three frontier models: GPT-5.2, Claude Sonnet 4.5, and Gemini 2.5 Pro.
+As shown in Table 2, the performance advantage of C2C remains consistent across all backbones.
+The Claude Sonnet 4.5 model, known for its specialized coding capabilities, achieves the highest raw performance.
+However, without constraints, it still incurs significant communication overhead in the Free Communication baseline (Efficiency 1.96).
+Applying the C2C framework boosts its efficiency to 2.40 (a 22% relative improvement).
+Similarly, GPT-5.2 sees an efficiency gain from 2.08 to 2.22, and Gemini 2.5 Pro from 1.88 to 2.09.
+These results confirm that while stronger models can execute tasks faster, they are not immune to coordination friction.
+C2C provides a universal optimization layer that scales with model capability, ensuring that high performance agents focus on execution rather than redundant communication.
+Backbone Model Policy Time (h) $\downarrow$ Comm Cost (h) $\downarrow$ Efficiency $\uparrow$
+GPT-5.2 Free Comm $11.50$ $4.20$ $2.08$
+C2C (Ours) $\mathbf{10.80}$ $\mathbf{2.15}$ $\mathbf{2.22}$
+Claude Sonnet 4.5 Free Comm $12.25$ $4.80$ $1.96$
+C2C (Ours) $\mathbf{10.00}$ $\mathbf{2.40}$ $\mathbf{2.40}$
+Gemini 2.5 Pro Free Comm $12.75$ $5.10$ $1.88$
+C2C (Ours) $\mathbf{11.50}$ $\mathbf{2.65}$ $\mathbf{2.09}$
+Table 2: Generalization analysis on Medium complexity tasks.
+Even with highly capable models, the unconstrained Free Comm baseline suffers from communication overhead.
+C2C consistently improves Efficiency by minimizing coordination costs, demonstrating that the framework’s value scales with model capability.
+4.4 Scalability and Multi-Task Analysis
+[…]
 Configuration Completion Time Comm/Agent Comm Cost Speedup
@@
 1M+8W (2 tasks) 21h 2.3 4.64 1.35
-Table 2: Scalability and multi-task analysis showing communication cost and speedup across team configurations.
-Multi-task performance demonstrates effective workload balancing and resource sharing capabilities.
-Speedup is calculated as the ratio of the completion time under the "No Communication" baseline to the completion time of each respective configuration.
-4.3 Scalability and Multi-Task Analysis
-Table 2 demonstrates C2C’s effectiveness across team sizes and workload scenarios.
-The framework exhibits sub-linear scaling in communication cost: while team size increases to 3.4 times (5 to 17 agents), communication cost only increases 86%.
-This favorable scaling behavior stems from C2C’s intelligent communication routing and the alignment factor mechanism, which prioritizes high value interactions over communications.
-Speedup analysis reveals that performance steadily improves as the number of agents increases, with the largest configuration tested (1M+16W) achieving the highest speedup of $\mathbf{1.95\times}$.
-This highlights the effective task parallelization and specialized skill utilization in larger teams.
-However, the incremental gain in speedup lessens as the team grows, suggesting that simply adding more workers yields progressively smaller advantages.
+Table 3: Scalability and multi-task analysis. Speedup is relative to the "No Communication" baseline.
+[…]
 The multi-task evaluation with 1M+8W reveals C2C’s ability to handle concurrent workloads effectively.
@@
 Message Type Analysis
-Table 3 details the impact of different message types on collaboration.
-MEETING_INVITE messages provide the highest alignment gains (+0.27), followed by HELP_REQUEST (+0.15), directly correlating with task success by resolving critical blockers.
-In contrast, PROGRESS_UPDATE messages maintain synchronization with a neutral (0) impact on alignment in this context.
-This analysis validates C2C’s strategy of prioritizing communications that address specific knowledge gaps.
+Table 4 confirms that MEETING_INVITE messages provide the highest alignment gains (+0.27), justifying their high cost for critical synchronization.
+HELP_REQUEST follows (+0.15), while PROGRESS_UPDATE has negligible immediate impact on AF but maintains long-term context.
 Message Type Frequency Avg Response Step Impact on Alignment
@@
 RESPONSE 8
-Table 3: Analysis of message types in the 1M+8W configuration on Medium tasks, showing the frequency and impact of each communication type.
+Table 4: Analysis of message types showing frequency and impact on AF.
 Task Decomposition Quality
-Effective collaboration begins with high quality task decomposition.
-We compare C2C’s adaptive decomposition strategy against manual, naive LLM, and hierarchical methods in Table 4.
-C2C’s approach, which considers agent skills and workload during decomposition, achieves superior subtask clarity and higher AF.
-This leads to 14% increase in worker utilization rate compared to naive LLM decomposition, demonstrating that context aware planning is critical for multi-agent efficiency.
+As shown in Table 5, C2C’s adaptive decomposition strategy achieves higher subtask clarity (0.95) compared to naive LLM prompting (0.72), approaching human manual quality (1.00).
 Decomposition Method Subtask Clarity Alignment Factor
@@
 C2C Adaptive 0.95 0.68
-Table 4: Comparison of task decomposition quality. Subtask Clarity is rated by an LLM judge (1.0 is best). C2C’s adaptive method approaches manual quality.
+Table 5: Comparison of task decomposition quality judged by LLM.
 4.6 Communication Pattern Analysis
@@
 5 Conclusion
-In this paper, we present Communication to Completion, a multi-agent framework that treats communication as an optimizable resource and quantifies task understanding through the Alignment Factor within a Sequential Action Framework.
-By treating alignment as a quantifiable and dynamic variable, C2C enables agents to autonomously determine when and how to communicate.
-Our experiments demonstrate that this alignment-driven approach adapts naturally to task complexity, with agents communicating more frequently and through richer channels when facing difficult problems while minimizing overhead on simpler tasks.
-Across diverse task complexities and team sizes, C2C consistently achieves high task completion while reducing completion time and improving efficiency relative to baselines.
-The framework produces interpretable patterns, including alignment factor trajectories, communication type distributions, and collaboration network structures, which bridge agent behavior and coordination theory to provide design insights for multi-agent systems.
+In this paper, we investigate the modeling of collaboration efficiency and communication constraints in multi-agent systems.
+Rather than focusing solely on agent architecture, we propose a multi—agent framework, Communication to Completion (C2C), to quantify the interplay between information alignment and coordination costs.
+By formalizing task understanding through the Alignment Factor and modeling communication as a scarce resource with explicit latency, we establish a rigorous methodology for quantifying the trade-offs between coordination overhead and execution efficiency in collaborative workplace.
+Our experiments validate this modeling approach, revealing that efficiency is not merely a function of individual agent capability, but a dynamic equilibrium between alignment gains and communication overhead.
+The simulation results demonstrate that imposing realistic constraints drives agents to emerge with strategic behaviors, such as batching questions or escalating to meetings, which are absent in unconstrained environments.
+These patterns hold consistent across different backbone models , suggesting that our cost-alignment model captures fundamental dynamics of collaborative intelligence.
 Limitations
-While the C2C framework demonstrates significant performance gains, we acknowledge several limitations that offer avenues for future research.
-The findings are derived from a controlled simulation, and the framework’s performance in unpredictable, real-world environments is yet to be validated.
-The experiments were confined to the software engineering domain, so the C2C’s effectiveness may not generalize to other collaborative fields.
-Furthermore, the Alignment Factor (AF) relies on an LLM’s evaluation, which introduces potential subjectivity and bias.
+While the C2C framework demonstrates significant efficiency gains, we acknowledge several limitations that offer avenues for future research.
+First, the Alignment Factor acts as a proxy for understanding but does not guarantee functional correctness (e.g., passing unit tests).
+Second, our deterministic cost models simplify real world stochastic interruptions.
+Finally, our experiments are specific to software engineering; generalizing these dynamics to other collaborative domains remains future work.
 Appendix A Tasks Details